OpenStudy (anonymous):

how do i write this hyperbola equation y^2-4x^2=1 without the 4 in front of the x so that i can graph it?

OpenStudy (anonymous):

Don't you want y all alone to be able to graph it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Right, so let's try moving stuff so y is all that's left on one side.

OpenStudy (anonymous):

y^2=4x^2+1

OpenStudy (anonymous):

Looks good

OpenStudy (anonymous):

What are you thinking now?

OpenStudy (anonymous):

Oh wait, i'm suppose to graph it in the x2-y2=1 form.

OpenStudy (anonymous):

so i need to get rid of the 4 here: y^2-4x^2=1

OpenStudy (anonymous):

You just want to sketch the graph right?

OpenStudy (anonymous):

I don't think you're supposed to remove the coefficient 4, but rather understand what effect it has on the graph

OpenStudy (anonymous):

My teacher wants me to do all my hw a certain way. Oh, I realize that I can put 1/4 under the x^2. Thanks anyway.

OpenStudy (anonymous):

:P If you do 1/4 under x^2, you have to do that to all other terms as well

OpenStudy (anonymous):

(y^2)/4 -x^2 = 1/4

OpenStudy (anonymous):

 I'm doing it this way?